OpenStudy (anonymous):

Find integral (9t+5)^2.6 dt

OpenStudy (anonymous):

(9t+5)^2.7 dt

OpenStudy (anonymous):

u = 9t+5 du = 9dt dt= du/9 integrate (1/9) * u^(2.7) *du

OpenStudy (anonymous):

use this rule: \[\int\limits_{ }^{ } x^n = \frac{ x^{n+1} }{ n+1 }\]

OpenStudy (anonymous):

i got my answer, but not sure if it is right. can u check @Algebraic!

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

1/33 u^3.7

OpenStudy (anonymous):

what i did is using substitution, u = 9t+5 du = 9 dt 1/9 du = dt